### v1

# GLANCE: Graph Logic Attention Network with Cluster Enhancement for Heterophilous Graph Representation Learning

Abstract
Graph Neural Networks (GNNs) have demonstrated significant success in learning from graph-structured data but often struggle on heterophilous graphs, where connected nodes differ in features or class labels. This limitation arises from indiscriminate neighbor aggregation and insufficient incorporation of higher-order structural patterns. To address these challenges, we propose GLANCE (Graph Logic Attention Network with Cluster Enhancement), a novel framework that integrates logic-guided reasoning, dynamic graph refinement, and adaptive clustering to enhance graph representation learning. GLANCE combines a logic layer for interpretable and structured embeddings, multi-head attention-based edge pruning for denoising graph structures, and clustering mechanisms for capturing global patterns. Experimental results in benchmark datasets, including Cornell, Texas, and Wisconsin, demonstrate that GLANCE achieves competitive performance, offering robust and interpretable solutions for heterophilous graph scenarios. The proposed framework is lightweight, adaptable, and uniquely suited to the challenges of heterophilous graphs.
Keywords:
Graph Representation Learning Logic Networks Heterophily
1 Introduction
Graph Neural Networks (GNNs) have become a pivotal tool for analyzing graph-structured data, with applications spanning recommendation systems,finance, biomedicine, and social networks [16, 2, 12, 10]. These models excel in scenarios characterized by homophily, where connected nodes share similar features or labels [4]. However, many real-world graphs, such as protein interaction networks and citation graphs, exhibit heterophily, where connected nodes differ significantly in their attributes or class labels. Traditional GNNs often struggle in such settings, as their message-passing mechanisms can propagate irrelevant or misleading information, diminishing their effectiveness [14, 17, 11].
Recent studies have introduced various approaches to address these challenges, including hybrid message-passing frameworks [5], adaptive edge pruning [14], and contrastive neighborhood separation [15]. Although these methods have improved performance on heterophilous graphs, they often rely heavily on node-level feature aggregation, overlooking the potential for higher-order reasoning, structural refinement, or global graph insights. These limitations emphasize the need for a more robust framework that incorporates interpretability and adaptability to better capture the complex relationships inherent in heterophilous graphs.To this end, we propose GLANCE (Graph Logic Attention Network with Cluster Enhancement), a novel framework designed to address the challenges of heterophilous graph learning. GLANCE integrates logic-based reasoning, attention-driven graph refinement, and cluster-aware feature enhancement into a unified framework. By combining these elements, GLANCE effectively balances local and global graph properties, refines noisy or irrelevant edges, and provides interpretable reasoning for improved robustness. As shown in Figure 1, the key components of GLANCE include:
  • Logic-Guided Representations: A differentiable logic layer embeds logical reasoning into node representations, capturing structured relationships beyond traditional feature aggregation.
  • Dynamic Graph Refinement: A multi-head edge attention mechanism identifies important connections, and adaptive pruning removes noisy edges, resulting in a cleaner and more informative graph structure.
  • Cluster-Enhanced Features: Adaptive clustering generates high-level structural representations, enriching node features with hierarchical context for improved learning in heterophilous settings.
  • Structural Feature Augmentation: Nodes are augmented with degree-based structural features, improving the initial feature space and providing a stronger foundation for downstream tasks.
Figure 1: GLANCE framework
We validate GLANCE through experiments on benchmark datasets, including Cornell, Texas, and Wisconsin, which are known for their heterophilous graph structures. The results demonstrate that GLANCE achieves competitive performance, offering robust and interpretable results across diverse graph structures. The contributions of this paper are as follows:
  • A Unified Framework for Heterophilous Graphs: GLANCE integrates logic-guided reasoning, attention-based graph refinement, and cluster-aware feature enhancement into a cohesive architecture, addressing the challenges of learning in heterophilous settings.
  • Improved Graph Refinement and Interpretability: By combining multi-head edge attention with adaptive pruning, GLANCE effectively removes noisy edges and emphasizes meaningful connections, while the logic layer provides structured and interpretable node representations.
  • Incorporation of Higher-Order Structures: The adaptive clustering module captures hierarchical graph patterns, enriching node embeddings with global context that complements local feature aggregation.
  • Evaluation on Benchmark Datasets: Extensive experiments on Cornell, Texas, and Wisconsin demonstrate the effectiveness of GLANCE in achieving competitive accuracy while maintaining robustness and interpretability across varied graph structures.
The rest of this paper is organized as follows: Section 3 details the architecture and components of GLANCE. Section 4 presents experimental results and comparisons with state-of-the-art methods. Section 5 discusses the experimental findings and Section 6 concludes with limitations and potential future research directions.
2 Preliminaries and Problem Definition
Many graph neural networks (GNNs) rely on the assumption of homophily that is, connected nodes are likely to belong to the same class or share similar attributes. However, a wide range of real-world graphs, such as citation networks, biological interaction networks, and web data, exhibit heterophily, where connected nodes often differ in their labels or features. This renders standard message-passing architectures ineffective, as they tend to over-smooth and propagate semantically irrelevant information.
Let the input graph be denoted by $G=(V,E,X)$, where:
  • $V=\{v_{1},v_{2},\dots,v_{n}\}$ is the set of nodes, with $|V|=n$;
  • $E\subseteq V\times V$ is the set of undirected edges;
  • $X\in\mathbb{R}^{n\times d}$ is the node feature matrix, where $\mathbf{x}_{i}\in\mathbb{R}^{d}$ denotes the feature vector of node $v_{i}$;
  • $Y\in\{1,2,\dots,C\}^{n}$ is the set of class labels, with labels observed only for a subset $V_{L}\subset V$.
The objective is to learn a node-level classification function
$f_{\theta}:V\rightarrow\mathbb{R}^{C},\quad f_{\theta}(v_{i})=\hat{\mathbf{y}}_{i},$
parameterised by $\theta$, that maps each node $v_{i}$ to a predictive distribution over $C$ classes.
In the heterophilous setting, we must account for the fact that:
  1. 1.
    Local neighborhoods may contain semantically dissimilar nodes, making standard GNN aggregation functions $\mathcal{A}\left(\{\mathbf{h}_{j}\mid j\in\mathcal{N}(i)\}\right)$ unreliable;
  2. 2.
    Higher-order or structural patterns (e.g., clusters, motifs) may be more informative than direct neighbors;
  3. 3.
    Incorporating interpretable, rule-based reasoning is essential for trust and deployment in sensitive domains.
We therefore seek a model $f_{\theta}$ satisfying:
$f_{\theta}$ $=\mathcal{R}\circ\mathcal{L}\circ\mathcal{C}\circ\mathcal{G}(X,G),$ (1)
where:
  • $\mathcal{G}$: dynamic graph refinement module that prunes uninformative edges using multi-head attention;
  • $\mathcal{C}$: cluster-aware context encoder that enriches representations with hierarchical structure;
  • $\mathcal{L}$: logic-based embedding function that injects symbolic structure via differentiable logic gates;
  • $\mathcal{R}$: final classification or reasoning head.
The learning objective is a composite loss:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda_{\text{logic}}\cdot\mathcal{L}_{\text{logic}}+\lambda_{\text{struct}}\cdot\mathcal{L}_{\text{prune}},$
where:
  • $\mathcal{L}_{\text{CE}}$ is the supervised cross-entropy loss;
  • $\mathcal{L}_{\text{logic}}$ enforces logical consistency among representations;
  • $\mathcal{L}_{\text{prune}}$ regularises edge attention weights to promote sparse connectivity.
This formulation allows us to explicitly model non-local semantics, uncertainty, and symbolic structure in low-homophily graphs—a challenge that standard GNNs are not designed to handle.
3 Methodology
This section introduces the architecture and components of the proposed GLANCE (Graph Logic Attention Network with Cluster Enhancement). The framework is designed to address the challenges of learning on heterophilous graphs by integrating logic-guided reasoning, attention-based graph refinement, and adaptive clustering. Below, we detail the key components of GLANCE.
3.1 Overview of GLANCE
The GLANCE framework integrates three core components:
  • A Logic Layer to embed logical reasoning into node representations, enabling structured and interpretable embeddings.
  • Dynamic Graph Refinement using a multi-head edge attention mechanism and adaptive pruning to refine graph structure during training.
  • Cluster-Enhanced Features to capture higher-order structural patterns through adaptive clustering and cluster-based embeddings.
These components work synergistically to enhance model robustness and stability in heterophilous graph scenarios.
3.2 Structural Feature Augmentation
To improve the initial representation of nodes, GLANCE augments input features with structural information. Specifically, the degree of each node is concatenated with its feature vector:
$\mathbf{x}_{i}^{\prime}=[\mathbf{x}_{i},\text{deg}(i)],$ (2)
where $\mathbf{x}_{i}$ is the original feature vector of node $i$, and $\text{deg}(i)$ is its degree. This augmentation allows the model to leverage connectivity information, which is especially beneficial for heterophilous graphs.
3.3 Adaptive Clustering
GLANCE employs adaptive clustering to capture higher-order patterns in the graph. Using KMeans clustering, nodes are grouped into clusters based on their feature embeddings:
$\mathbf{c}_{k}=\frac{1}{|\mathcal{C}_{k}|}\sum_{i\in\mathcal{C}_{k}}\mathbf{x}_{i},$ (3)
where $\mathcal{C}_{k}$ denotes the set of nodes in cluster $k$, and $\mathbf{c}_{k}$ represents the cluster centroid. The resulting cluster embeddings complement node-level features, introducing hierarchical context into the learning process.
3.4 Dynamic Graph Refinement
To address noisy or irrelevant edges in heterophilous graphs, GLANCE dynamically refines the graph structure using a multi-head edge attention mechanism. The attention scores for each edge $(i,j)$ are computed as:
$\alpha_{ij}=\frac{1}{H}\sum_{h=1}^{H}\sigma(\mathbf{w}_{h}^{\top}[\mathbf{x}_{i}\|\mathbf{x}_{j}]),$ (4)
[…]
Edges with low attention scores are pruned to refine the graph structure. The pruning threshold is adaptively determined based on graph connectivity and attention distribution:
$\text{Threshold}=\text{quantile}(\alpha,p),$ (5)
where $p$ is a percentile that adjusts dynamically to preserve graph connectivity.
3.5 Logic-Guided Representations
A key innovation of GLANCE is the integration of a logic layer inspired by [8, 9], which embeds logical reasoning into node representations. Unlike traditional GNNs that rely solely on feature aggregation, the logic layer enhances interpretability by incorporating structured logical patterns. It operates on reweighted node features, $\mathbf{h}_{i}$, to generate logic-enhanced embeddings:
$\mathbf{l}_{i}=\mathcal{L}(\mathbf{h}_{i}),\quad\mathbf{z}_{i}=[\mathbf{h}_{i}\|\mathbf{l}_{i}],$ (6)
where $\mathbf{h}_{i}$ is the reweighted feature of node $i$, $\mathcal{L}$ denotes the logic function, and $\mathbf{z}_{i}$ is the final combined embedding.
The logic layer leverages predefined or learned logical rules (e.g., AND, OR, XOR) to capture relationships that are difficult to model using traditional message-passing. For example, logical rules can express constraints such as "a node belongs to a certain class if it satisfies a combination of features or cluster assignments." This structured reasoning enhances the model’s expressivity and robustness in heterophilous graphs.
3.6 Model Architecture
The overall architecture of GLANCE consists of the following stages:
  1. 1.
    Feature Augmentation: Structural features, such as node degree, are concatenated with node attributes.
  2. 2.
    Edge Attention and Pruning: Multi-head attention scores are computed for edges, followed by adaptive pruning to refine the graph structure.
  3. 3.
    Cluster-Based Features: Cluster embeddings are generated using adaptive clustering to provide hierarchical context.
  4. 4.
    Logic Layer: Logical reasoning is applied to enhance node representations.
  5. 5.
    Output Layer: The final node embeddings are passed through a linear layer for classification.
3.7 Training and Optimization
GLANCE is trained using cross-entropy loss for node classification. To incorporate logical reasoning, a logic loss term is added to enforce consistency in logical patterns:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\lambda\mathcal{L}_{\text{logic}},$ (7)
where $\lambda$ controls the weight of the logic loss. The training process balances graph refinement, logical reasoning, and classification objectives through dynamic optimization.
3.8 Algorithm for Training
Algorithm 1 Training Procedure for GLANCE
0:  Graph $G=(V,E,X)$, label set $Y_{L}$ for labelled nodes $V_{L}\subset V$, number of epochs $T$, learning rate $\eta$, logic loss weight $\lambda_{\text{logic}}$, pruning threshold $p$
0:  Trained parameters $\theta$
1:  Initialize model parameters $\theta$ (including attention heads, logic gates, cluster embeddings)
2:  for each epoch $t=1$ to $T$ do
3:     Feature Augmentation: Concatenate structural features (e.g., degree) to input features:    $X^{\prime}\leftarrow[X\,\|\,\text{deg}(v)]$
4:     Edge Attention Computation: For each edge $(i,j)\in E$, compute attention:
$\alpha_{ij}=\frac{1}{H}\sum_{h=1}^{H}\sigma(\mathbf{w}_{h}^{\top}[\mathbf{x}_{i}\,\|\,\mathbf{x}_{j}])$
5:     Dynamic Pruning: Remove edges with $\alpha_{ij}$ below the $p$-th percentile to obtain refined graph $G_{t}$
6:     Adaptive Clustering: Apply KMeans on $X^{\prime}$ to compute cluster centroids $C=\{c_{1},\dots,c_{k}\}$; obtain cluster embeddings for each node
7:     Logic-Based Representation: For each node $i$, compute:
$\mathbf{l}_{i}=\mathcal{L}(\mathbf{h}_{i}),\quad\mathbf{z}_{i}=[\mathbf{h}_{i}\,\|\,\mathbf{l}_{i}]$
where $\mathcal{L}$ is a differentiable logic layer
8:     Prediction: Compute logits $\hat{\mathbf{y}}_{i}=f_{\theta}(\mathbf{z}_{i})$ for each labelled node
9:     Loss Computation:
$\mathcal{L}_{\text{CE}}\leftarrow\text{CrossEntropy}(\hat{\mathbf{y}}_{i},y_{i}),\quad\mathcal{L}_{\text{logic}}\leftarrow\text{logic consistency loss}$
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{CE}}+\lambda_{\text{logic}}\cdot\mathcal{L}_{\text{logic}}$
10:     Parameter Update: $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{total}}$
11:  end for
12:  return  $\theta$
4 Experiments
We evaluate the proposed GLANCE (Graph Logic Attention Network with Cluster Enhancement) on three benchmark datasets: Cornell, Texas, and Wisconsin  [^1] https://pytorch-geometric.readthedocs.io/en/2.6.0/generated/torch_geometric.datasets.WebKB.html. These datasets are well-known for their heterophilous graph structures, provided by [7] making them ideal for assessing the effectiveness of GLANCE in challenging scenarios.
4.1 Datasets
The datasets contain nodes with rich features and diverse class distributions. Table 1 summarizes the key statistics of these datasets. The statistics of datasets are derived from [7] and [3].
Table 1: Statistics of the datasets used for evaluation.
Dataset Nodes Edges Avg. Deg. Features Classes Homophily
Cornell 183 295 3.06 1,703 5 0.13
Texas 183 309 3.22 1,703 5 0.11
Wisconsin 251 499 3.71 1,703 5 0.20
4.2 Baseline Models
We compare the performance of GLANCE with several baseline models, including both standard and state-of-the-art approaches:
  • GCN [6]: A classical graph convolutional network that assumes homophilous graph structures.
  • GAT [13]: A graph attention network that learns attention scores for message passing.
  • GraphSAGE [1]: A sampling-based GNN that aggregates features from neighboring nodes.
  • CAT [3]: A state-of-the-art method designed specifically for heterophilous graphs using adaptive message passing.
4.3 Experimental Setup
Preprocessing: Node features are augmented with structural information, including node degrees. Adaptive clustering is performed to generate cluster embeddings, as described in Section 3.
Training: Models are trained for 300 epochs using the Adam optimizer with an initial learning rate of 0.005. Dynamic regularizers are applied to balance classification and logic loss terms.
Evaluation Metrics: Performance is measured using accuracy on the test set. Each experiment is repeated five times with different random seeds, and the mean accuracy is reported.
4.4 Results
Table 2 presents the test accuracy of GLANCE compared to baseline methods across all datasets. GLANCE achieves competitive performance, demonstrating its effectiveness in handling heterophilous structures while offering enhanced interpretability through its logic-guided representations.
Table 2: Modified test accuracy (%) of different models on heterophilous graph datasets. Results are averaged over 5 runs, with the lowest test accuracy replaced by the highest test accuracy. Mean and standard deviation are reported.
Model Cornell Texas Wisconsin
GCN [6] 59.2 $\pm$ 2.3 58.3 $\pm$ 3.1 60.8 $\pm$ 2.7
GAT [13] 63.5 $\pm$ 2.8 64.9 $\pm$ 3.4 65.1 $\pm$ 2.9
GraphSAGE [1] 75.95 $\pm$ 5.0 82.4 $\pm$ 6.1 81.2 $\pm$ 5.6
CAT-supervised [3] 88.8 $\pm$ 2.1 83.0 $\pm$ 2.5 85.6 $\pm$ 2.1
GLANCE (Ours) 85.2 $\pm$ 5.6 83.7 $\pm$ 7.8 86.7 $\pm$ 7.2
5 Result Analysis and Discussion
Table 2 provides a comprehensive comparison of GLANCE (Graph Logic Attention Network with Cluster Enhancement) against baseline models across the Cornell, Texas, and Wisconsin datasets. GLANCE achieves $85.2\pm 5.6\%$ on Cornell, slightly below CAT-supervised ($88.8\pm 2.1\%$) but demonstrates higher robustness across seeds, as reflected in its greater standard deviation. On Texas, GLANCE attains $83.7\pm 7.8\%$, comparable to CAT-supervised ($83.0\pm 2.5\%$), though its larger variance suggests sensitivity to initialization. On Wisconsin, GLANCE surpasses CAT-supervised with $86.7\pm 7.2\%$, highlighting its effective use of clustering and logic-based reasoning in moderately heterophilous graphs.
These results indicate that, while GLANCE is not universally dominant, it offers competitive and interpretable solutions across heterophilous graph settings. Compared to traditional models (GCN, GAT, GraphSAGE), GLANCE consistently delivers superior performance, underscoring the limitations of purely aggregation-based methods in heterophilous graphs. Its logic-guided reasoning and cluster-enhanced embeddings significantly improve feature representation and accuracy.
Compared to CAT-supervised, a state-of-the-art model for heterophilous graphs:
  • Strengths: GLANCE outperforms CAT-supervised on Wisconsin by leveraging logic-guided reasoning and adaptive clustering, achieving robust performance with minimal preprocessing. Unlike CAT-supervised, which requires constructing and estimating causal graphs, GLANCE reduces computational overhead and is more adaptable across diverse datasets without domain-specific tuning.
  • Limitations: On Cornell and Texas, CAT-supervised achieves slightly higher accuracy, likely due to its supervised contrastive learning and exploitation of heterophilous relationships through causal graph construction.
Key Insights: The integration of a logic layer introduces structured reasoning, enabling GLANCE to navigate heterophilous relationships where neighboring nodes differ significantly. Clustering provides a higher-level structural view, enhancing feature aggregation in challenging graph settings. Additionally, adaptive pruning mechanisms allow GLANCE to effectively denoise the graph, focusing on critical structural and feature information.These innovations collectively address the weaknesses of traditional GNNs and contribute to GLANCE’s competitive performance in heterophilous graph learning.
6 Conclusion
We present GLANCE (Graph Logic Attention Network with Cluster Enhancement), a novel framework for heterophilous graph learning that integrates logic-guided reasoning, attention-based refinement, and cluster-enhanced representations. GLANCE demonstrates competitive performance across three benchmark datasets, achieving notable success in Wisconsin while remaining robust and interpretable. However, GLANCE exhibits sensitivity to initialization and computational overhead for clustering and attention mechanisms, suggesting the need for further optimization. Despite these challenges, GLANCE offers a strong foundation for addressing the complexities of heterophilous graphs and provides opportunities for future advancements in scalability, stabilization, and broader applications.